In the summer of 2010, Mr.Butler purchased a pre-owned Ford Expedition that averaged 12.1 miles per gallon. In June 2013 the transmission needed to be rebuild at a cost of $3200. After the transmission was rebuilt the average gas mileage increased to 13.4 miles per gallon. Assuming the cost of gas is $3.50 per gallon, how many miles must be driven for the gas savings to be the same as the cost of the transmission?
Answer by TimothyLamb(4379)

fuel economy improvement = 13.4 - 12.1 = 1.3 miles/gallon
---
(1/1.3 gallons/mile) * (x miles) * (3.50 $/gallon) = 3200
(1/1.3) * x * 3.50 = 3200
x = 3200/((1/1.3) * 3.50)
x = 1188.56 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php